OpenStudy (anonymous):

find any point (a,b) on the curve x^2+xy=2y-8 where the tangent line at the point is vertical?

OpenStudy (anonymous):

take the derivative and find where it is undefined

OpenStudy (anonymous):

\[x^2+xy=2y-8\] \[2x+y+xy'=2y'\] solve for \(y'\)

OpenStudy (anonymous):

\[2x+y=2y'-xy'\] \[2x+y=(2-x)y'\] \[\frac{2x+y}{2-x}=y'\] undefined if \(x=2\)